OpenStudy (anonymous):

If the simple interest on $1000 for 5 years is $250, then what is the rate?

OpenStudy (legends):

@Nurali Is it 4%

OpenStudy (legends):

?

OpenStudy (legends):

@Jamie1019 Is there any answer choices

OpenStudy (anonymous):

no

OpenStudy (legends):

Its 5% rate

OpenStudy (nurali):

I = Prt

OpenStudy (nurali):

I=250 P=1000 t=5

OpenStudy (nurali):

r=I/Pt

OpenStudy (legends):

0.05*1,000*5=250

OpenStudy (nurali):

r=250/1000*5 r=0.05 r=0.05*100 r=5%

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (legends):

I was right!! WOHOO

OpenStudy (legends):

@Jamie1019 Are you there